# Testing Dependency of Weighted Random Graphs

### v1

@@ Abstract
In this paper, we study the task of detecting the edge dependency between two weighted random graphs. We formulate this task as a simple hypothesis testing problem, where under the null hypothesis, the two observed graphs are statistically independent, whereas under the alternative, the edges of one graph are dependent on the edges of a randomly vertex-permuted version of the other graph. For general edge-weights distributions, we establish thresholds at which optimal testing is information-theoretically impossible and possible, as a function of the total number of nodes in the observed graphs and the generative distributions of the weights. Finally, we observe a statistical-computational gap in our problem, and we provide evidence that this is fundamental using the framework of low-degree polynomials.
@@ 1 Introduction
Consider the following decision problem. We observe two weighted random graphs that are either generated independently at random or are edge-dependent under some latent vertex correspondence/permutation. For this basic problem, two natural questions arise: the detection problem, i.e., determining whether the graphs are dependent or not, and the recovery problem, i.e., identifying the underlying vertex correspondence. Here, we address the former question, specifically, we aim to understand under what conditions, in terms of the number of vertices and the generative distributions, one can distinguish between the two hypotheses and detect whether these graphs are dependent or not, say, with high probability?
The fundamental question above was first introduced and analyzed in [1], where for Gaussian-weighted and dense Erdős-Rényi random graphs on $n$ vertices, sharp information-theoretic thresholds were developed, revealing the exact barrier at which the asymptotic optimal detection error probability undergoes a phase transition from zero to one as $n$ approaches infinity. For sparse Erdős-Rényi random graphs this threshold was initially determined within a constant factor in the same paper. Later on, this result was refined in [2], where a sharp information-theoretic threshold was established for a specific range of edge probabilities. Our work is closely related to these works, and much inspired by [1]; we study the detection limit under general weight distributions, beyond the Erdős-Rényi and Gaussian Wigner graphs models above, and thereby aim to answer [1, Sec. 6, Problem 2].
@@
Recently, there has been a growing interest in the problem of random graph matching, also known as network alignment problem on the detection of two random graphs, and more generally, two databases, as well as the closely related problem of matching the vertex correspondence in the presence of correlation. This study is greatly motivated by many applications in applied fields, such as, computational biology applications [4, 5], social network analysis and data anonymization/privacy-focused systems [6, 7], computer vision [8, 9], just to name a few. In addition to [1], which is most related to our paper, we mention a few other related works briefly. The information-theoretic thresholds for both the problems of detection and matching (recovery) are fairly well-understood under the Erdős-Rényi and Gaussian Wigner probabilistic models. For detection, significant progress has been made in establishing sharp thresholds and understanding the fundamental limits of exact recovery, as shown in works such as [10, 11, 1, 12]. For matching and recovery, advancements include achieving polynomial-time algorithms for graph similarity, identifying information-theoretic thresholds for vertex matching, and developing methods for partial recovery and alignment, as demonstrated in [13, 14, 15, 16, 17]. Furthermore, from the algorithmic point of view, a verity of polynomial-time efficient algorithms have been proposed along with rigorous performance guarantees. These include approaches for graph alignment and matching, such as those based on message-passing techniques, likelihood ratio tests, and spectral methods. Notable advances include algorithms for exact recovery in correlated Erdős-Rényi graphs graphs, matching in Gaussian Wigner models, and partial alignment in sparse random graphs, as demonstrated in [14, 18, 19, 20, 21, 22, 23, 3, 24]. These algorithms have been essential in providing practical solutions to the graph matching problem, although they often fall short of the theoretical optimum. Interestingly, current belief is that there is a statistical-computational gap in the detection and recovery problems, in the sense that state-of-the-art efficient algorithms thresholds are strictly larger than the information-theoretic ones, and as so there is a statistical cost for computational efficiency. Quite recently, an evidence for this gap was given in [25] for Erdős-Rényi graphs based on the low-degree polynomial conjecture. It was demonstrated that low-degree polynomial algorithms fail for detection in both dense and sparse regimes, with specific results showing that any degree-$d$ polynomial algorithm fails for detection when the edge correlation is above a certain threshold, and in the sparse regime, when the edge density is sufficiently low. This suggests that several state-of-the-art algorithms for correlation detection and exact matching recovery may be close to optimal, reinforcing the statistical-computational gap. Finally, we mention here that the network alignment problem is closely related to a wide verity of other planted matching problems, both conceptually and technically, e.g., the database alignment and detection problem, with many exciting and interesting results, see, e.g., [26, 27, 28, 29, 30, 31], and many references therein.
@@ 2 Problem Formulation
Inspired by the hypothesis testing model proposed by [1], we formulate a general problem of testing network dependency as follows. Let $G=([n],\mathsf{W})$ denote a weighted undirected graph on the node set $[n]$ with weighted adjacency matrix $\mathsf{W}$, where $\mathsf{W}_{ii}=0$, for any $1\leq i<j\leq n$, and $\mathsf{W}_{ij}=1$ (or the edge weight) if $i$ and $j$ are adjacent and $\mathsf{W}_{ij}=0$, otherwise. Recall that two weighted graphs $G=([n],\mathsf{W})$ and $H=([n],\mathsf{W}^{\prime})$ are isomorphic and denoted by $G\cong H$ if there exists a permutation (called a graph isomorphism) $\pi$ on $[n]$ such that $\mathsf{W}_{ij}=\mathsf{W}^{\prime}_{\pi_{i}\pi_{j}}$, for all $i,j$. Given a weighted graph $G$, its isomorphism class ${\cal G}$ is the equivalence class ${\cal G}=\{H:H\cong G\}$. We refer to an isomorphism class as an unlabeled graph and ${\cal G}$ as the unlabeled version of $G$.
Problem 1.
Let $G_{1}=([n],\mathsf{W})$ and $G_{2}=([n],\mathsf{W}^{\prime})$ be two weighted random graphs, where the edge weights $\{(\mathsf{W}_{ij},\mathsf{W}_{ij}^{\prime}):1\leq i<j\leq n\}$ are i.i.d. pairs of random variables, and $\mathsf{W}_{ij}$ and $\mathsf{W}_{ij}^{\prime}$ have the same marginal distribution. Under the null hypothesis ${\cal H}_{0}$, $\mathsf{W}_{ij}$ and $\mathsf{W}_{ij}^{\prime}$ are independent; under the alternative hypothesis ${\cal H}_{1}$, $\mathsf{W}_{ij}$ and $\mathsf{W}_{ij}^{\prime}$ are correlated. Given the unlabeled versions of $G_{1}$ and $G_{2}$, i.e., their isomorphism classes ${\cal G}_{1}$ and ${\cal G}_{2}$, the goal is to test ${\cal H}_{0}$ versus ${\cal H}_{1}$.
We now reformulate the testing problem given in Problem 1 in a more convenient form. Due to the exchangeability of the (i.i.d.) edge weights, observing the unlabeled version is equivalent to observing its randomly relabeled version. Indeed, let $\pi$ and $\pi^{\prime}$ be two independent random permutations uniformly drawn from the set $\mathbb{S}_{n}$. Consider the relabeled version of $G_{1}=([n],\mathsf{W})$ with weighted adjacency matrix $\mathsf{A}$, where $\mathsf{A}_{ij}=\mathsf{W}_{\pi_{i}\pi_{j}}$; similarly, let $\mathsf{B}$ correspond to the relabeled version of $G_{2}=([n],\mathsf{W}^{\prime})$ with $\mathsf{B}_{ij}=\mathsf{W}^{\prime}_{\pi^{\prime}_{i}\pi^{\prime}_{j}}$. It is clear that observing the unlabeled graphs ${\cal G}_{1}$ and ${\cal G}_{2}$ is equivalent to observing the labeled graphs $\mathsf{A}$ and $\mathsf{B}$. Since $\pi^{\prime-1}\circ\pi$ is also a uniform random permutation, we arrive at the following formulation that is equivalent to Problem 1.
Problem 2.
Let $G=([n],\mathsf{A})$ and $G^{\prime}=([n],\mathsf{B})$ denote weighted undirected graphs on the vertex set $[n]$, with weighted adjacency matrices $\mathsf{A}$ and $\mathsf{B}$, where $\mathsf{A}_{ii}=0$, for any $1\leq i<j\leq n$, and $\mathsf{A}_{ij}=1$ (or the edge weight) if $i$ and $j$ are adjacent and $\mathsf{A}_{ij}=0$, otherwise, and similarly for $\mathsf{B}$. We assume that the edge weights of each graph (equivalently, $\{\mathsf{A}_{ij}\}_{1\leq i<j\leq n}$ and $\{\mathsf{B}_{ij}\}_{1\leq i<j\leq n}$) are i.i.d. Now, under the null hypothesis ${\cal H}_{0}$, the matrices $\mathsf{A}$ and $\mathsf{B}$ are statistically independent, with $\mathsf{A}_{i,j}\sim{\cal P}_{{A}}$ and $\mathsf{B}_{i,j}\sim{\cal P}_{{B}}$, for all $i,j\in[n]$, respectively. It is assumed that ${\cal P}_{{A}}={\cal P}_{{B}}$, and we denote the product measure ${\cal Q}_{{A}{B}}={\cal P}_{{A}}\times{\cal P}_{{B}}$. Under the alternative hypothesis ${\cal H}_{1}$, given a latent random permutation $\pi\sim\mathsf{Unif}(\mathbb{S}_{n})$, the random variables $\{(\mathsf{A}_{ij},\mathsf{B}_{\pi_{i}\pi_{j}}):1\leq i<j\leq n\}$ are i.i.d., and each pair $(\mathsf{A}_{ij},\mathsf{B}_{\pi_{i}\pi_{j}})\sim{\cal P}_{{A}{B}}$, with the same marginals as ${\cal Q}_{{A}{B}}$, i.e., ${\cal P}_{{A}}={\cal P}_{{B}}$. Upon observing $\mathsf{A}$ and $\mathsf{B}$, the goal is to test ${\cal H}_{0}$ versus ${\cal H}_{1}$. For clarity and brevity, we will denote univariate random variables by $A$ and $B$ and adjacency matrices by $\mathsf{A}$ and $\mathsf{B}$ throughout this paper.
To summarize, the hypothesis testing problem we deal with is,
@@
A test function for our problem is a map $\phi:\{0,1\}^{n\times n}\times\{0,1\}^{n\times n}\to\{0,1\}$, designed to determine which hypothesis ${\cal H}_{0}$ or ${\cal H}_{1}$ occurred. We define the risk of a test $\phi$ as the sum of its Type-I and II error probabilities, i.e.,
@@
We remark that $\mathsf{R}$ is a function of $n$, however, we omit them from our notation for the benefit of readability. We study the possibility and impossibility of our detection problem in multiple asymptotic regimes. We say that $\phi$ achieves strong detection if $\mathsf{R}_{n}(\phi)\to 0$, as $n\to\infty$, and weak detection if $\limsup_{n\to\infty}\mathsf{R}_{n}(\phi)<1$ (i.e., strictly outperforming a random guess). Our goal is to investigate when strong/weak detection is possible and impossible, as a function of $n$ and the distributions $({\cal P},{\cal Q})$.
@@ 4.1.2 A formula for the second moment
Importantly, observe that $X_{O}$ is a function of $(\mathsf{A}_{ij},\mathsf{B}_{\pi_{i},\pi_{j}})_{(i,j)\in O}$. Let ${\cal O}$ denote the collection of all edge orbits of $\sigma^{\mathsf{E}}$. Since edge orbits are disjoint, we have
@@
Since $\{\mathsf{A}_{ij}\}_{i<j}$ and $\{\mathsf{B}_{ij}\}_{i<j}$ are independent under ${\cal H}_{0}$, we conclude $\{\mathsf{X_{O}}\}_{O\in{\cal O}}$ are mutually independent under ${\cal H}_{0}$. Therefore,

### v2

@@ Abstract
In this paper, we study the task of detecting the edge dependency between two weighted random graphs. We formulate this task as a simple hypothesis testing problem, where under the null hypothesis, the two observed graphs are statistically independent, whereas under the alternative, the edges of one graph are dependent on the edges of a uniformly and randomly vertex-permuted version of the other graph. For general edge-weight distributions, we establish thresholds at which optimal testing becomes information-theoretically possible or impossible, as a function of the total number of nodes in the observed graphs and the generative distributions of the weights. Finally, we identify a statistical-computational gap, and present evidence suggesting that this gap is inherent using the framework of low-degree polynomials.
@@ 1 Introduction
Consider the following decision problem. We observe two weighted random graphs that are either generated independently at random or are edge-dependent due to some latent vertex correspondence or permutation. For this basic problem, two natural questions arise: the detection problem, which concerns whether the graphs exhibit dependence, and the recovery problem, which concerns identifying the latent correspondence between vertices. Here, we address the former question, specifically, we aim to understand under what conditions, in terms of the number of vertices and the generative distributions, one can distinguish between the two hypotheses and detect whether these graphs are dependent or not, say, with high probability?
The fundamental question above was first introduced and analyzed in [1], where for Gaussian-weighted and dense Erdős-Rényi random graphs on $n$ vertices, sharp information-theoretic thresholds were developed, revealing the exact barrier at which the asymptotic optimal detection error probability undergoes a phase transition from zero to one as $n$ approaches infinity. For sparse Erdős-Rényi random graphs this threshold was initially determined within a constant factor in the same paper. Later on, this result was refined in [2], where a sharp information-theoretic threshold was established for a specific range of edge probabilities. Our work is closely related to these works, and much inspired by [1]; we study the detection limit under general weight distributions, beyond the Erdős-Rényi and Gaussian Wigner graphs models mentioned above, and thereby aim to answer [1, Sec. 6, Problem 2].
@@
Recently, there has been a growing interest in the problem of random graph matching, also known as network alignment problem on the detection of two random graphs, and more generally, two databases, as well as the closely related problem of matching the vertex correspondence in the presence of correlation. This research is driven by a broad spectrum of practical applications, covering areas such as computational biology [4, 5], social network analysis, data anonymization and privacy-focused systems [6, 7], and computer vision [8, 9], among others. In addition to [1], which is most related to our paper, we mention a few other related works briefly. The information-theoretic thresholds for both the problems of detection and matching (recovery) are fairly well-understood under the Erdős-Rényi and Gaussian Wigner probabilistic models. For detection, significant progress has been made in establishing sharp thresholds and understanding the fundamental limits of exact recovery, as shown in works such as [10, 11, 1, 12]. For matching and recovery, advancements include achieving polynomial-time algorithms for graph similarity, identifying information-theoretic thresholds for vertex matching, and developing methods for partial recovery and alignment, as demonstrated in [13, 14, 15, 16, 17]. Furthermore, from an algorithmic perspective, a variety of polynomial-time efficient algorithms have been proposed along with rigorous performance guarantees. These include approaches for graph alignment and matching, such as those based on message-passing techniques, likelihood ratio tests, and spectral methods. Notable advances include algorithms for exact recovery in correlated Erdős-Rényi graphs graphs, matching in Gaussian Wigner models, and partial alignment in sparse random graphs, as demonstrated in [14, 18, 19, 20, 21, 22, 23, 3, 24]. These algorithms have been essential in providing practical solutions to the graph matching problem, although they often fall short of the theoretical optimum. Interestingly, current belief is that there is a statistical-computational gap in the detection and recovery problems, in the sense that state-of-the-art efficient algorithms thresholds are strictly larger than the information-theoretic ones, and as so there is a statistical cost for computational efficiency. Quite recently, an evidence for this gap was given in [25] for Erdős-Rényi graphs based on the low-degree polynomial conjecture. It was demonstrated that low-degree polynomial algorithms fail for detection in both dense and sparse regimes, with specific results showing that any degree-$d$ polynomial algorithm fails for detection when the edge correlation is above a certain threshold, and in the sparse regime, when the edge density is sufficiently low. This suggests that several state-of-the-art algorithms for correlation detection and exact matching recovery may be close to optimal, reinforcing the statistical-computational gap. Finally, we mention here that the network alignment problem is closely related to a wide variety of other planted matching problems, both conceptually and technically, e.g., the database alignment and detection problem, with many exciting and interesting results, see, e.g., [26, 27, 28, 29, 30, 31], and many references therein.
@@ 2 Problem Formulation
We follow the formulation of testing problem proposed in [1], but generalize to arbitrary weighted random graphs. Specifically, let $G=([n],\mathsf{A})$ and $G^{\prime}=([n],\mathsf{B})$ denote weighted undirected graphs on the vertex set $[n]$, with weighted adjacency matrices $\mathsf{A}\in\mathbb{R}^{n\times n}$ and $\mathsf{B}\in\mathbb{R}^{n\times n}$, respectively. We assume that the edge weights of each graph, namely, $\{\mathsf{A}_{ij}\}_{1\leq i<j\leq n}$ and $\{\mathsf{B}_{ij}\}_{1\leq i<j\leq n}$, to be i.i.d. Under the null hypothesis ${\cal H}_{0}$, the matrices $\mathsf{A}$ and $\mathsf{B}$ are statistically independent, with $\mathsf{A}_{i,j}\sim{\cal P}_{{A}}$ and $\mathsf{B}_{i,j}\sim{\cal P}_{{B}}$, for all $i,j\in[n]$, respectively. It is assumed that ${\cal P}_{{A}}={\cal P}_{{B}}$, and we denote the product measure ${\cal Q}_{{A}{B}}={\cal P}_{{A}}\times{\cal P}_{{B}}$. Under the alternative hypothesis ${\cal H}_{1}$, given a latent random permutation $\pi\sim\mathsf{Unif}(\mathbb{S}_{n})$, the random variables $\{(\mathsf{A}_{ij},\mathsf{B}_{\pi_{i}\pi_{j}}):1\leq i<j\leq n\}$ are i.i.d., and each pair $(\mathsf{A}_{ij},\mathsf{B}_{\pi_{i}\pi_{j}})\sim{\cal P}_{{A}{B}}$, with the same marginals as ${\cal Q}_{{A}{B}}$, i.e., ${\cal P}_{{A}}={\cal P}_{{B}}$. Upon observing $\mathsf{A}$ and $\mathsf{B}$, the objective is to evaluate ${\cal H}_{0}$ against ${\cal H}_{1}$. For clarity and brevity, we will denote univariate random variables by $A$ and $B$ and adjacency matrices by $\mathsf{A}$ and $\mathsf{B}$ throughout this paper. To summarize, the hypothesis testing problem we deal with is,
@@
A test function for our problem is a map $\phi:\{0,1\}^{n\times n}\times\{0,1\}^{n\times n}\to\{0,1\}$, designed to determine which hypothesis ${\cal H}_{0}$ or ${\cal H}_{1}$ occurred. ”We characterize the risk associated with a test $\phi$ the total of its Type-I and II error probabilities, i.e.,
@@
We remark that $\mathsf{R}$ is a function of $n$, however, we omit them from our notation for the benefit of readability. We study the possibility and impossibility of our detection problem in multiple asymptotic regimes. A test $\phi$ achieves strong detection if $\mathsf{R}_{n}(\phi)$ approaches 00 as $n$ tends to infinity, and weak detection if $\limsup_{n\to\infty}\mathsf{R}_{n}(\phi)<1$ (i.e., strictly outperforming a random guess). Our main goal is to investigate when strong/weak detection is possible and impossible, as a function of $n$ and the distributions $({\cal P},{\cal Q})$.
@@ 4.1.2 A formula for the second moment
Next, we note that $X_{O}$ is a function of $(\mathsf{A}_{ij},\mathsf{B}_{\pi_{i},\pi_{j}})_{(i,j)\in O}$. Accordingly, we let ${\cal O}$ denote the partition of edge orbits of $\sigma^{\mathsf{E}}$. Since edge orbits are disjoint, we have
@@
Since $\{\mathsf{A}_{ij}\}_{i<j}$ and $\{\mathsf{B}_{ij}\}_{i<j}$ are independent under ${\cal H}_{0}$, we conclude $\{\mathsf{X_{O}}\}_{O\in{\cal O}}$ are statistically independent under ${\cal H}_{0}$. Therefore,
@@ 5 Conclusion and Outlook
In this paper, we analyzed the problem of detecting the dependency weighted random graphs. To that end, we derived both lower and upper bounds on the phase transition boundaries at which detection is possible or impossible. We show that these bounds are tight in several special cases considered in the literature, and in other distributions as well. We observed a gap between the statistical limits we derive and the performance of the efficient algorithms we construct. We conjecture that this gap is in fact inherent, and to provide an evidence for this conjecture we follow show that the class of low-degree polynomials fail the conjectureally hard regime. We hope our work has opened more doors than it closes. Let us mention briefly a few interesting questions for future research:
  1. 1.
    It would be interesting to devise universal algorithms which are independent of the underlying generative distributions, which assumed known in our paper.
  2. 2.
    In this paper, we investigated the average-case/noisy version of the famous worst-case graph isomorphism problem; accordingly it would be interesting and important to analyze the related subgraph isomorphism problem, where only a subgraph (say, chosen uniformly at random) in the original graph is correlated with a node-permuted version in the other graph.
  3. 3.
    Similarly to [20], it is quite important to devise efficient algorithms that achieve strong detection (rather than weak detection as we proposed in our paper).